OpenStudy (anonymous):

solve by completeing the square 2x^2-7x-17+

OpenStudy (anonymous):

2x^2-7x-17=0

OpenStudy (anonymous):

Completing the square, for any quadratic equation:\[ax^2 +bx + c=0\] 1. divide both sides by a 2. subtract c from both sides 3. add\[\frac{b^2}{4}\]to both sides 4. rewrite the left-hand side as a perfect square 5. square root both sides 6. solve for x.